Margin of Safety

Comer Company produces and sells strings of colorful indoor/outdoor lights for holiday display to retailers for $9.52 per string. The variable costs per string are as follows:

Direct materials $1.87
Direct labor 1.70
Variable factory overhead 0.57
Variable selling expense 0.42

Fixed manufacturing cost totals $297,600 per year. Administrative cost (all fixed) totals $225,680. Comer expects to sell 241,500 strings of light next year.

Required:

1. Calculate the break-even point in units.
units

2. Calculate the margin of safety in units.
units

3. Calculate the margin of safety in dollars.
$
Answer #1

Fixed cost= 297,600+225,680 = 523,280

Contribution margin per unit = Sales - Variable costs

= 9.52 - (1.87+1.70+0.57+0.42) = 4.96

1)

Breakeven point in units

= Fixed cost/Contribution margin per unit

= 523,280/4.96

= 105,500

2)

Margin of Safety = Sales - Breakeven sales

= 241,500 - 105,500

= 136,000

​​​​​​3)

Margin of Safety = 136,000*9.52

= 1,294,720